OpenStudy (anonymous):

By decreasing each side of a rectangle by 1 unit, the area decreased from 60 square feet to 44 square feet. Find the perfect decrease in area.

OpenStudy (dumbcow):

L*W = 60 (L-1)(W-1) = 44 --> LW -L -W +1 = 44 By substitution L+W = 17 L = 17 - W W(17-W) = 60 W^2 -17W +60 = 0 (W-12)(W-5) = 0 W = 5 or 12 Dimensions of original rectangle are 12ft by 5ft

OpenStudy (anonymous):

But I don't see how that has to do with the percent decrease of the area of the two rectangles?

OpenStudy (xkehaulanix):

This one I'm not completely sure about, but here goes. Since you start out with 60 sq ft and end with 44 sq ft, you want to find out the percent lost from 60. Start out with how much was lost: 60 - 44 = 16 Now 16 sq ft was the amount lost from 60 sq ft. What percentage is 16 from 60? 16/60 = .2666666666666666666666... ...which rounds to about .2667 To get the percentage, multiply .2667 by 100% .2667 x 100% = 26.67% Percentages definitely were not my strong suit in math, so take this with a grain of salt.

OpenStudy (anonymous):

That was right! thank you!